In a shipment of 100 televisions, 6 are defective. If a person buys two televisions from that shipment, what is the probability that both are defective?
Answer by edjones(8007)

6/100 chance 1st one is defective.
5/99 chance 2nd is defective
6/100 * 5/99 = 30/990
=1/33
.
Another way:
nCr=2C6=15
nCr=2C100=4950
15/4950=1/33
.
Ed